Problem 3: Using a Spreadsheet to Calculate Pension Benefit Payments

Your employer uses a final pay formula to determine retirement payments to its employees. You have 20 years of service at the company and are considering retirement sometime in the next 10 years. Your employer uses a final pay formula by which you receive an annual benefit payment of 4 percent of your employer uses a final pay formula by which you receive an annual benefit payment of 4 percent of your average salary over the last three years of service times the number of years employed. Calculate the annual benefit if you retire now, in 2 years, 5 years, 8 years, and 10 years using the estimated salary during the last three years of service listed below.

Average Years of Benefit => Annual  
Retire Salary Service Percent Payment
now $50,000 ENTER VALUE = FORMULA
in 2 years $51,005 =
in 5 years $52,551 =
in 8 years $54,143 =
in 10 years $55,231 =
Answer #1

The only tricky part here is to calculate the last 3 year average salary. This is clear for 5 years and 8 years because for the last three years, we can consider the given average salary as the salary for 3 years. However, for 2 years and 10 years, we will ahve to calculate the average salary for last three years from the given information. The details of the calculation is copied in the screenshot below:

If this understanding of using last 3 year salary is incorrect then we will have to take the given average salary as given.

A В с Е F G Н 1 Comment Average Years of Annual 2 Benefit => 3 Retire Salary Service Percent Payment Avg for last 3 years con

The excel with formula copied below:

Average Years of Benefit => Annual   Comment
Retire Salary Service Percent Payment
now 50000 20 =0.04*D4 = =E4*C4 Avg for last 3 years considered as $50,000
in 2 years 51005 22 =0.04*D5 = =(C5*2+C4)/3*E5 Avg for last 3 years calculated as ($50,000 + 2*$51005)/3 because we need to consider average of last 3 years
in 5 years 52551 25 =0.04*D6 = =C6*E6 Avg for last 3 years considered as $52,551
in 8 years 54143 28 =0.04*D7 = =C7*E7 Avg for last 3 years considered as $54,143
in 10 years 55231 30 =0.04*D8 = =(C8*2+C7)/3*E8 Avg for last 3 years calculated as ($54,143 + 2*$55,231)/3 because we need to consider average of last 3 years